OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. x to the 1 third power times x to the 1 third power x to the 1 third power x to the 2 sixths power x to the 2 thirds power x to the 1 ninth power

OpenStudy (anonymous):



OpenStudy (dan815):

\[x^n*x^k=x^{n+k}\]

OpenStudy (johnweldon1993):

\[\large x^{1/3} \times x^{1/3} = x^{1/3 + 1/3} = x^{2/3}\] You just add the exponents

OpenStudy (anonymous):

ohhh C

OpenStudy (anonymous):

@dan815 @johnweldon1993 thank you both so muchhh